Calculate the mass, in tonnes, of calcium carbonate that is required to react (in an aqueous...

  1. Calculate the mass, in tonnes, of calcium carbonate that is required to react (in an aqueous environment) with one tonne of carbon dioxide (which becomes chemically trapped as calcium bicarbonate, Ca(HCO3)2). (Note: one tonne equals 1,000 kg.)
Answer #1

CaCO + Con+ H₂O - 1 1 1 Ca(HCO3) moar 1000 kg 0 Yox= 100/kg = 1000 x103 g / 4 g/mol 37 g/mol No= 1000kg = 1000 x 10 g 44 g/mo